# inkscapelatex=false

### v1

@@ 6. Symmetric value matrix
7. Simulations
@@
In this section we illustrate the theoretical results and show that its conclusions appear to hold even when our assumptions are violated. First, we simulate the continuous transformer model and provide an example for each of our theoretical results, i.e., for Theorems 3.24.15.1, and  6.1. Then, we use the GPT-2 XL model to illustrate how token consensus seems to occur even if the assumptions in our theoretical results are not satisfied. These experiments also illustrate the effects of token consensus on the decoding process.
In Sections 7.1 to 7.3 we use $10$ three-dimensional tokens with random initial positions. The initial position of the tokens is represented by a white circle and final position by a gray circle. Their trajectories are displayed in blue. All matrices are initially chosen randomly, with each element drawn from a uniform distribution on the interval $[-0.5,0.5]$. When a specific assumption needs to be satisfied, we modify the randomly chosen matrix to construct one that adheres to the required properties. For example, if we assume $P$ is symmetric, we start with a random matrix $A$ and compute $P$ as $P=A+A^{\top}$.
7.1. Illustration of Theorem 3.2
On the left of Figure 1 we can appreciate the motion of 10 tokens on the ellipsoid $(\mathcal{E}_{P}^{2})^{10}$ defined by the randomly generated positive definite symmetric matrix:
$P=\begin{pmatrix}0.9432&0.0587&-0.1813\\ 0.0587&0.8450&-0.1013\\ -0.1813&-0.1013&0.5519\end{pmatrix},$
according to the dynamics (8). As expected, all of the tokens converge to a consensus equilibrium. In this case, the dynamics is a gradient vector field and on the right of Figure 1 we show the time evolution of the corresponding potential $V_{P}=V\circ\imath_{P}$ with $V$ defined in (10).
(a) Convergence to a consensus equilibrium on the ellipsoid $\mathcal{E}^{2}_{P}$.
(b) Evolution of the potential function $V_{P}=V\circ\imath_{P}$ with $V$ defined in (10).
Figure 1. Illustration of Theorem 3.2.
7.2. Illustration of Theorem 4.1
@@
Figure 3 shows the motion of 10 tokens on the sphere $(\mathbb{S}^{2})^{10}$ according to the dynamics (11) with $h=2$. The matrices $P_{1}(t)$ and $P_{2}(t)$ were computed as $P_{1}(t)=D_{1}(t)P_{1}^{\prime}$, $P_{2}(t)=D_{2}(t)P_{2}^{\prime}$ with $P_{1}^{\prime}$ and $P_{2}^{\prime}$ randomly generated:
$P_{1}^{\prime}=\begin{pmatrix}0.0805&-0.1929&0.1991\\ -0.2312&0.3131&-0.2335\\ 0.1788&-0.1732&-0.1594\end{pmatrix},\quad P_{2}^{\prime}=\begin{pmatrix}-0.3067&0.0349&0.1107\\ 0.0572&-0.0557&0.1343\\ 0.1375&0.1083&0.1018\end{pmatrix},$
and $D_{1}(t)$ and $D_{2}(t)$ given by:
$D_{1}=\begin{pmatrix}2\cos(10\pi t)&0&0\\ 0&2\sin(10\pi t)&0\\ 0&0&2\cos(6\pi t)\end{pmatrix},\quad D_{2}=\begin{pmatrix}2\cos(6\pi t)&0&0\\ 0&2\sin(6\pi t)&0\\ 0&0&2\cos(4\pi t)\end{pmatrix}.$
To better appreciate the time-varying nature of the matrices $P_{1}$ and $P_{2}$, in Figure 2 we shown their Frobenius norm.
Figure 2. Frobenius norm of the matrices $P_{1}$ and $P_{2}$.
@@
(a) Convergence to a consensus equilibrium on the sphere $\mathbb{S}^{2}$. All the tokens start and remain in an hemisphere.
(b) Evolution of the Lyapunov function (13) used in the proof of Theorem 4.1.
Figure 3. Illustration of Theorem 4.1.
@@
On the left of Figure 3 we can appreciate that all the tokens start and remain in an hemisphere and that they converge to a consensus equilibrium. The proof of Theorem 4.1 was based on the Lyapunov function (13) whose time-evolution is displayed on the right of Figure 3 for the case where $v=(1~{}0~{}0)$.
7.3. Illustration of Theorem 6.1
Theorem 6.1 considers the dynamics (21) with $P(t)=D(t)P^{\prime}$, randomly generated $P^{\prime}$ and $U$ given by:
$P^{\prime}=\begin{pmatrix}0.3598&0.4150&0.1319\\ 0.0971&-0.0668&-0.2046\\ 0.1548&-0.2102&0.1220\end{pmatrix},\quad U=\begin{pmatrix}-0.2590&0.4965&0.5609\\ 0.4965&-0.7174&-0.5003\\ 0.5609&-0.5003&-0.0247\end{pmatrix},$
and $D(t)$ defined as:
$D=\begin{pmatrix}2\cos(10\pi t)&0&0\\ 0&2\sin(10\pi t)&0\\ 0&0&2\cos(6\pi t)\end{pmatrix}.$
On the left of Figure 4 we can observe convergence of the tokens to a consensus equilibrium point whereas on the right we have the time evolution of $V_{1}=1-{\rm y}_{1}^{\top}v$ and $V_{2}=1-{\rm y}_{2}^{\top}v$ where $v\in\mathbb{R}^{3}$ is the eigenvector of $U$ corresponding to its largest eigenvalue. Note that $V_{2}$ is not a Lyapunov function, and therefore it may increase, although the proof of Theorem 6.1, establishes that it will eventually converge to zero.
(a) Convergence to a consensus equilibrium on the sphere $\mathbb{S}^{2}$. All the tokens start and remain in the hemisphere defined by $v$.
(b) Evolution of the functions $V_{1}$ and $V_{2}$ used in the proof of Theorem 6.1.
Figure 4. Illustration of Theorem 6.1.
7.4. Illustration of Theorem 5.1
@@
We now consider the dynamics (15) with $50$ tokens each of dimension $500$. The number and dimension of the tokens were chosen to make them comparable to the GPT-2 model. We use two heads ($h=2$) with the matrices $P_{1}=D_{1}(t)P_{1}^{\prime}$ and $P_{2}=D_{2}(t)P_{2}^{\prime}$ obtained by randomly generating $P_{1}^{\prime}$ and $P_{2}^{\prime}$, and taking $D_{1}(t)$ and $D_{2}(t)$ to be diagonal with entries $(D_{\eta})_{jj}=|2\sin(wt+\phi)|$ for $\eta=1,2$, $j=1,\ldots,500$, $w$ drawn from the uniform distribution on ]0,1[]0,1[] 0 , 1 [ and $\phi$ drawn from the uniform distribution on $]0,2\pi[$.
In Figure 5 we display the evolution of the function $E:(\mathcal{E}_{W}^{n})^{\ell}\to\mathbb{R}^{+}$ defined as:
(25) $E=1-\frac{1}{\ell}\sum_{i=0}^{\ell}\frac{|{\rm y}_{1}^{\top}\,{\rm y}_{i}|}{|{\rm y}_{1}|\,|{\rm y}_{i}|},$
@@
along $100$ trajectories of (15) from random initial conditions drawn from an element-wise uniform distribution on ]0.5,0.5[]-0.5,0.5[] - 0.5 , 0.5 [, and then projected to the sphere. The function $E$ only becomes zero when all the tokens are located at the same configuration. We can appreciate in Figure 5 how the function $E$ converges to zero along all the trajectories.
@@
Figure 5. Illustration of Theorem 5.1; evolution of the function $E$ defined in (25) along $100$ solutions of (15) with random initial conditions drawn from an element-wise uniform distribution on ]0.5,0.5[]-0.5,0.5[] - 0.5 , 0.5 [ and then projected to the sphere.
7.5. GPT-2 Experiments

### v2

@@ 5.2. Symmetric value matrix
5.3. Time-varying value matrix
Lastly, we extend the result in the previous section to time-varying value matrices. In this case, the eigenvectors corresponding to the maximum eigenvalue of the value matrix are also time-varying. Our result states that the tokens converge to some neighborhood of this time-varying eigenvalue if they start on the hemisphere defined by the eigenvalue at the initial time.
Hypothesis 5.3.
The model Eq.˜6 is auto-regressive, $W=\mathbb{I}_{n+1}$, and we have:
  1. (1)
    There is only one head, i.e., $h=1$,
  2. (2)
    $U_{1}(t)=U(t)$ is differentiable (as a function of $t$), bounded and symmetric, i.e., $\sup_{t\in\mathbb{R}_{0}^{+}}\|U(t)\|<\infty$ and $U^{\top}(t)=U(t)$, and
  3. (3)
    $P_{1}(t)=P(t)$ is bounded, i.e., $\sup_{t\in\mathbb{R}_{0}^{+}}\|P(t)\|<\infty$.
For each $t\in\mathbb{R}_{0}^{+}$, we denote the spectrum of $U(t)$ by $\lambda(U(t))=\{\lambda_{1}(t),\ldots,\lambda_{r}(t)\}$ with $\lambda_{\mu}(t)>\lambda_{\nu}(t)$ for each $1\leq\mu<\nu\leq r$. Note that $\lambda(U(t))\subset\mathbb{R}$ as $U(t)$ is symmetric. Given $\lambda(t)\in\lambda(U(t))$, the corresponding eigenspace is denoted by $L_{\lambda(t)}(U(t))\subset\mathbb{R}^{n+1}$.
Let us introduce a family of functions that will be useful for the next result. An almost class $\mathcal{K}\mathcal{L}$ function is continuous function $\Theta:\mathbb{R}_{0}^{+}\times\mathbb{R}_{0}^{+}\to\mathbb{R}_{0}^{+}$ such that, for each $(r_{0},t_{0})\in\mathbb{R}_{0}^{+}\times\mathbb{R}_{0}^{+}$, we have:
  1. (1)
    the function $\mathbb{R}_{0}^{+}\ni r\mapsto\Theta(r,t_{0})\in\mathbb{R}_{0}^{+}$ is strictly increasing, and
  2. (2)
    there exists $\alpha\in\mathcal{K}$ such that $\lim_{t\to\infty}\Theta(r_{0},t)=\alpha(r_{0})$.
Theorem 5.3.
For each $t\in\mathbb{R}_{0}^{+}$, let $v_{1}(t)\in L_{\lambda_{1}(t)}(U(t))\cap\mathbb{S}^{n}$ and define $\varepsilon=\sup_{t\in\mathbb{R}_{0}^{+}}\{|\dot{v}_{1}(t)|\}$. If ˜5.3 holds and:
  1. (1)
    $\lambda_{1}(t)>0$ for each $t\in\mathbb{R}_{0}^{+}$,
  2. (2)
    $\dim L_{\lambda_{1}(t)}(U(t))=1$ for each $t\in\mathbb{R}_{0}^{+}$, and
  3. (3)
    $\delta=\inf_{t\in\mathbb{R}_{0}^{+}}(\lambda_{1}(t)-\lambda_{2}(t))>0$,
then, for each $1\leq i\leq\ell$, there exist:
  1. (1)
    a class $\mathcal{KL}$ function $\beta_{i}:\mathbb{R}_{0}^{+}\times\mathbb{R}_{0}^{+}\to\mathbb{R}_{0}^{+}$, and
  2. (2)
    an almost class $\mathcal{K}\mathcal{L}$ function $\Theta_{i}:\mathbb{R}_{0}^{+}\times\mathbb{R}_{0}^{+}\to\mathbb{R}_{0}^{+}$,,
such that, for each solution $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell}):\mathbb{R}_{0}^{+}\to(\mathbb{S}^{n})^{\ell}$ of Eq.˜6 with $\mathbf{y}(0)\in\mathcal{D}^{\ell}(v_{1}(0))$, we have:
$|{\rm y}_{i}(t)-v_{1}(t))|^{2}\leq\beta_{i}(|{\rm y}_{i}(0)-v_{i}(0)|^{2},t)+\Theta_{i}\left(\varepsilon,t\right),\quad 1\leq i\leq\ell,~t\in\mathbb{R}_{0}^{+}.$
Proof.
Firstly, note that Item˜2 of ˜5.3 and Item˜2 in the Theorem statement ensure that $\dot{v}_{1}:\mathbb{R}_{0}^{+}\to\mathbb{R}^{n+1}$ is defined almost everywhere. For each $t\in\mathbb{R}_{0}^{+}$, let $\{v_{1}(t),\ldots,v_{n+1}(t)\}$ be an orthonormal basis of eigenvectors of $U(t)$, which exists due to its symmetry. Let us denote by $\gamma_{\mu}(t)\in\mathbb{R}$ the eigenvalue corresponding to $v_{\mu}(t)$ for $1\leq\mu\leq n+1$. Note that $\gamma_{1}(t)=\lambda_{1}(t)$.
For each $1\leq i\leq\ell$, let $e_{i}={\rm y}_{i}-v_{1}:\mathbb{R}_{0}^{+}\to\mathbb{R}^{n+1}$ and $b_{i}=v_{1}^{\top}{\rm y}_{i}:\mathbb{R}_{0}^{+}\to[-1,1]$. Henceforth, we drop the argument $t$ for simplicity. By writing ${\rm y}_{i}=\sum_{\mu=1}^{n+1}(v_{\mu}^{\top}{\rm y}_{i})v_{\mu}$ we obtain $1=|{\rm y}_{i}|^{2}=\sum_{\mu=1}^{n+1}(v_{\mu}^{\top}{\rm y}_{i})^{2}$. This, as well as the fact that $U=\sum_{\mu=1}^{n+1}\gamma_{\mu}\,v_{\mu}v_{\mu}^{\top}$, yields:
$\lambda_{1}-{\rm y}_{i}^{\top}U{\rm y}_{i}$ $=\lambda_{1}-\sum_{\mu=1}^{n+1}\gamma_{\mu}(v_{\mu}^{\top}{\rm y}_{i})^{2}$
$=\lambda_{1}-\lambda_{1}\left(1-\sum_{\mu=2}^{n+1}(v_{\mu}^{\top}{\rm y}_{i})^{2}\right)-\sum_{\mu=2}^{n+1}\gamma_{\mu}(v_{\mu}^{\top}{\rm y}_{i})^{2}$
$=\sum_{\mu=2}^{n+1}(\lambda_{1}-\gamma_{\mu})(v_{\mu}^{\top}{\rm y}_{i})^{2}.$
From this and Eq.˜6 under the hypothesis in ˜5.3, the dynamics of $b_{1}$ reads:
$\dot{b}_{1}$ $=v_{1}^{\top}\dot{\rm y}_{1}+\dot{v}_{1}^{\top}{\rm y}_{1}$
$=\tilde{\alpha}_{11}\left(v_{1}^{\top}U{\rm y}_{1}-{\rm y}_{1}^{\top}U{\rm y}_{1}v_{1}^{\top}{\rm y}_{1}\right)+\dot{v}_{1}^{\top}{\rm y}_{1}$
$=\tilde{\alpha}_{11}\left(\lambda_{1}-{\rm y}_{1}^{\top}U{\rm y}_{1}\right)b_{1}+\dot{v}_{1}^{\top}{\rm y}_{1},$
(20) $=\tilde{\alpha}_{11}b_{1}\sum_{\mu=2}^{n+1}(\lambda_{1}-\gamma_{\mu})(v_{\mu}^{\top}{\rm y}_{1})^{2}+\dot{v}_{1}^{\top}{\rm y}_{1},$
for each $(t,b_{1})\in\mathbb{R}_{0}^{+}\times[-1,1]$. Similarly, the dynamics of $b_{i}$, $2\leq i\leq\ell$, reads:
$\dot{b}_{i}$ $=v_{1}^{\top}\dot{\rm y}_{i}+\dot{v}_{1}^{\top}{\rm y}_{i}$
$=\sum_{j=1}^{i-1}\tilde{\alpha}_{ij}(v_{1}^{\top}U(e_{j}+v_{1})-{\rm y}_{i}^{\top}U(e_{j}+v_{1})v_{1}^{\top}{\rm y}_{i})+\tilde{\alpha}_{ii}(v_{1}^{\top}U{\rm y}_{i}-{\rm y}_{i}^{\top}U{\rm y}_{i}v_{1}^{\top}{\rm y}_{i})+\dot{v}_{1}^{\top}{\rm y}_{i}$
$=\sum_{j=1}^{i-1}\tilde{\alpha}_{ij}(\lambda_{1}v_{1}^{\top}e_{j}+\lambda_{1}-({\rm y}_{i}^{\top}Ue_{j}+\lambda_{1}b_{i})b_{i})+\tilde{\alpha}_{ii}(\lambda_{1}-{\rm y}_{i}^{\top}U{\rm y}_{i})b_{i}+\dot{v}_{1}^{\top}{\rm y}_{i}$
$=\sum_{j=1}^{i-1}\tilde{\alpha}_{ij}\lambda_{1}(1-b_{i}^{2})+\tilde{\alpha}_{ii}(\lambda_{1}-{\rm y}_{i}^{\top}U{\rm y}_{i})b_{i}+g(e)+\dot{v}_{1}^{\top}{\rm y}_{i},$
(21) $=\sum_{j=1}^{i-1}\tilde{\alpha}_{ij}\lambda_{1}(1-b_{i}^{2})+\tilde{\alpha}_{ii}b_{i}\sum_{\mu=2}^{n+1}(\lambda_{1}-\gamma_{\mu})(v_{\mu}^{\top}{\rm y}_{i})^{2}+g(e)+\dot{v}_{1}^{\top}{\rm y}_{i},$
for each $(t,b_{i})\in\mathbb{R}_{0}^{+}\times[-1,1]$, where $e=(e_{1},\ldots,e_{i-1})$ and $g(e)=\sum_{j=1}^{i-1}\tilde{\alpha}_{ij}(\lambda_{1}v_{1}^{\top}-b_{i}{\rm y}_{i}^{\top}U)e_{j}$. Note that:
$|g(e)|$ $\leq\sum_{j=1}^{i-1}\tilde{\alpha}_{ij}(\lambda_{1}|v_{1}|+|b_{i}|\,|{\rm y}_{i}|\,\|U\|)|e_{j}|$
$\leq\sum_{j=1}^{i-1}c_{2}(\sup_{t\in\mathbb{R}_{0}^{+}}\lambda_{1}(t)+\sup_{t\in\mathbb{R}_{0}^{+}}\|U(t)\|)|e_{j}|$
(22) $=C\sum_{j=1}^{i-1}|e_{j}|,$
with $c_{2}>0$ as in Lemma˜4.1 and $C=c_{2}(\sup_{t\in\mathbb{R}_{0}^{+}}\lambda_{1}(t)+\sup_{t\in\mathbb{R}_{0}^{+}}\|U(t)\|)\in\mathbb{R}^{+}$, where we used Item˜2 and Item˜1 to ensure that $0<\sup_{t\in\mathbb{R}_{0}^{+}}\lambda_{1}(t)<\infty$.
Let us proceed by complete induction.
  1. (1)
    Base case. For $i=1$, Item˜2 and the orthogonality of the basis of eigenvectors chosen ensure that the only equilibria of Eq.˜20 are $b_{1}^{*}=-1$ (which corresponds to ${\rm y}_{1}^{*}=-v_{1}$) and $b_{1}^{*}=1$ (which corresponds to ${\rm y}_{1}^{*}=v_{1})$, when $\dot{v}_{1}=0$. In order to analyze the stability of the latter, let us consider the function $V_{+}:[-1,1]\to\mathbb{R}$ defined as $V_{+}(b_{1})=1-b_{1}$. Clearly, $V_{+}(1)=0$ and $V_{+}|_{]0,1[}>0$. Moreover, from Items˜2 and 3 and Eq.˜20, we obtain:
    $\dot{V}_{+}(t,b_{1})$ $=V_{+}^{\prime}(b_{1})\,\dot{b}_{1}\leq-\delta c_{1}(1-b_{1}^{2})b_{1}+\varepsilon,\qquad(t,b_{1})\in\mathbb{R}_{0}^{+}\times[0,1],$
    with $c_{1}>0$ as in Lemma˜4.1. For $\varepsilon=0$ and $t\in\mathbb{R}_{0}^{+}$, $\dot{V}_{+}(t,1)=0$ and $\dot{V}_{+}(t,\cdot)|_{]0,1[}<0$. This ensures that $V_{+}$ is a strict ISS-Lyapunov function for the equilibrium $b_{1}^{*}=1$ of Eq.˜20, where the input is given by $\dot{v}_{1}^{\top}{\rm y}_{1}$. We conclude that $b_{1}^{*}=1$ is ISS-stable with domain of attraction ]0,1[]0,1[. Hence, there exist $\tilde{\alpha}_{1}\in\mathcal{K}$ and $\tilde{\beta}_{1}\in\mathcal{KL}$ such that, for each solution $b_{1}:\mathbb{R}_{0}^{+}\to[-1,1]$ of Eq.˜20 with $b_{1}(0)\in]0,1]$ (which corresponds to ${\rm y}_{1}(0)\in\mathcal{D}(v_{1}(0))$), we have $|b_{1}(t)-1|\leq\tilde{\beta}_{1}(|b_{1}(0)-1|,t)+\tilde{\alpha}_{1}(\varepsilon)$ for each $t\in\mathbb{R}_{0}^{+}$. By noting that $|{\rm y}_{1}-v_{1}|=|e_{1}|=\sqrt{\langle e_{1},e_{1}\rangle}=\sqrt{2|b_{1}-1|}$, the previous condition may be rewritten as:
    $|{\rm y}_{1}(t)-v_{1}(t)|^{2}$ $=2|b_{1}(t)-1|$
    $\leq 2\tilde{\beta}_{1}(|b_{1}(0)-1|,t)+2\tilde{\alpha}_{1}(\varepsilon)$
    $=2\tilde{\beta}_{1}\left(\frac{|y_{1}(0)-v_{1}(0)|^{2}}{2},t\right)+2\tilde{\alpha}_{1}(\varepsilon),$
    for each $t\in\mathbb{R}_{0}^{+}$. By picking $\beta_{1}(r,t)=2\tilde{\beta}_{1}(r/2,t)$ and $\Theta_{1}(r,t)=2\tilde{\alpha}_{1}(r)$ for each $(r,t)\in\mathbb{R}_{0}^{+}\times\mathbb{R}_{0}^{+}$, we conclude.
  2. (2)
    Inductive step. Now, for $2\leq i\leq\ell$ let us assume that the result holds for the subsystem of Eq.˜6 under the assumptions in the statement given by the first i1i-1 tokens. Namely, there exist a class $\mathcal{KL}$ function $\beta_{j}:\mathbb{R}_{0}^{+}\times\mathbb{R}_{0}^{+}\to\mathbb{R}_{0}^{+}$ and an almost class $\mathcal{KL}$ function $\Theta_{j}:\mathbb{R}_{0}^{+}\times\mathbb{R}_{0}^{+}\to\mathbb{R}_{0}^{+}$, $1\leq j\leq i-1$, such that, for each solution $({\rm y}_{1},\ldots,{\rm y}_{j}):\mathbb{R}_{0}^{+}\to(\mathbb{S}^{n})^{i-1}$ of such subsystem with $({\rm y}_{1}(0),\ldots,{\rm y}_{j}(0))\in\mathcal{D}^{j}(v_{1}(0))$, we have:
    (23) $|e_{j}(t)|^{2}\leq\beta_{j}(|e_{j}(0)|^{2},t)+\Theta_{j}(\varepsilon,t),\qquad t\in\mathbb{R}_{0}^{+},~1\leq j\leq i-1.$
    Thanks to the orthogonality of the basis of eigenvectors, the only equilibria of Eq.˜21 are $b_{i}^{*}=-1$ (which corresponds to ${\rm y}_{i}^{*}=-v_{1}$) and $b_{i}^{*}=1$ (which corresponds to ${\rm y}_{i}^{*}=v_{1}$) when $e=(e_{1},\ldots,e_{i-1})=0$ and $\dot{v}_{1}=0$. To analyze the stability of the latter, let us consider the function $V_{+}:[-1,1]\to\mathbb{R}$ as defined above, i.e., $V_{+}(b_{i})=1-b_{i}$, which satisfies $V_{+}(1)=0$ and $V_{+}|_{]0,1[}>0$. From Items˜1, 2 and 3, Eq.˜21 and Eq.˜22, we obtain:
    (24) $\dot{V}_{+}(t,b_{i})$ $=V_{+}^{\prime}(b_{i})\,\dot{b}_{i}\leq-\delta c_{1}(1-b_{i}^{2})b_{i}+C\sum_{j=1}^{i-1}|e_{j}|+\varepsilon,$
    for each $(t,b_{i})\in\mathbb{R}_{0}^{+}\times[0,1]$. For $e=(e_{1},\ldots,e_{i-1})=0$ and $\varepsilon=0$, $\dot{V}_{+}(t,1)=0$ and $\dot{V}_{+}(t,\cdot)|_{]0,1[}<0$. Thus, $V_{+}$ is a strict ISS-Lyapunov function for the equilibrium $b_{i}^{*}=1$ of Eq.˜21, where the input is given by $\tilde{e}=C\sum_{j=1}^{i-1}|e_{j}|+\varepsilon$. Given that $b_{i}^{*}=1$ is ISS-stable with domain of attraction ]0,1[]0,1[, there exist $\tilde{\alpha}_{i}\in\mathcal{K}_{\infty}$ and $\tilde{\beta}_{i}\in\mathcal{KL}$ such that, for each solution $b_{i}:\mathbb{R}_{0}^{+}\to[-1,1]$ of Eq.˜21 with $b_{i}(0)\in]0,1]$ (which corresponds to ${\rm y}_{i}(0)\in\mathcal{D}(v_{1}(0))$) we have:
    $|b_{i}(t)-1|$ $\leq\tilde{\beta}_{i}(|b_{i}(0)-1|,t)+\tilde{\alpha}_{i}\left(C\sum_{j=1}^{i-1}|e_{j}|+\varepsilon\right),\qquad t\in\mathbb{R}_{0}^{+}.$
    From this, the fact that $|{\rm y}_{j}-v_{1}|=|e_{j}|=\sqrt{2|b_{j}-1|}$ for $1\leq j\leq i$ and Eq.˜23, we obtain:
    $|e_{i}(t)|^{2}=2|b_{i}(t)-1|\leq 2\tilde{\beta}_{i}(|b_{i}(0)-1|,t)+2\tilde{\alpha}_{i}\left(C\sum_{j=1}^{i-1}|e_{j}|+\varepsilon\right)$
    $\leq 2\tilde{\beta}_{i}\left(\frac{|e_{i}(0)|^{2}}{2},t\right)+2\tilde{\alpha}_{i}\left(C\sum_{j=1}^{i-1}\sqrt{\beta_{j}(2,t)+\Theta_{j}(\varepsilon,t)}+\varepsilon\right),$
    for each $t\in\mathbb{R}_{0}^{+}$, where we used that $|e_{j}(0)|\leq\sqrt{2}$ as ${\rm y}_{j}(0)\in\mathcal{D}(v_{1}(0))$, $1\leq j\leq i-1$. We conclude by choosing $\beta_{i}(r,t)=2\tilde{\beta}_{i}(r/2,t)$ and:
    $\Theta_{i}(r,t)=2\tilde{\alpha}_{i}\left(C\sum_{j=1}^{i-1}\sqrt{\beta_{j}(2,t)+\Theta_{j}(r,t)}+r\right).$
Under the conditions of Theorem˜5.3 , there exist $\alpha_{i}\in\mathcal{K}$ such that:
$\lim_{t\to\infty}|{\rm y}_{i}(t)-v_{1}(t)|^{2}\leq\alpha_{i}(\varepsilon),\qquad 1\leq i\leq\ell,$
provided $\mathbf{y}(0)=({\rm y}_{1}(0),\ldots,{\rm y}_{\ell}(0))\in\mathcal{D}^{\ell}(v_{1}(0))$. Therefore, all tokens converge to some ball around the time-varying consensus $\mathbf{y}^{*}(t)=(v_{1}(t),\ldots,v_{1}(t))$ as $t\to\infty$. The radius of the ball depends on $\varepsilon=\sup_{t\in\mathbb{R}_{0}^{+}}\{|\dot{v}_{1}(t)|\}$. Moreover, the basin of attraction is the open hemisphere generated by $v_{1}(0)$. For the limit case where $\varepsilon=0$, i.e., when $v_{1}(t)$ does not depend on time, we recover exact convergence to the consensus, as in Theorem˜5.2.
Remark 5.4 (Closest results available in the literature).
The authors are not aware of any previous results addressing the case where the value matrix is time-varying.
6. Simulations and empirical validation
@@
In this section we illustrate the theoretical results and show that their conclusions appear to hold even when our assumptions are violated. We start by simulating the continuous transformer model and illustrating our theoretical results. In addition to simulations, we provide empirical evidence using the GPT-2 XL and the GPT-Neo 2.7B to show how token consensus seems to occur even if the assumptions in our theoretical results are not satisfied.
6.1. Numerical simulations
6.1.1. Illustration of Theorem 4.1
@@ 6.1. Numerical simulations
We simulate the motion of 10 tokens, each of them randomly placed on the sphere $\mathbb{S}^{2}\subset\mathbb{R}^{3}$, according to the dynamics Eq.˜6 with $h=2$. All matrices, except for $P_{1}(t)$ and $P_{2}(t)$, were randomly chosen, and each element was drawn from a uniform distribution in the interval [0.5,0.5][-0.5,0.5]. The matrices $P_{1}(t)$ and $P_{2}(t)$ were computed as $P_{1}(t)=D_{1}(t)P_{1}^{\prime}$, $P_{2}(t)=D_{2}(t)P_{2}^{\prime}$ with $P_{1}^{\prime}$ and $P_{2}^{\prime}$ randomly generated:
$P_{1}^{\prime}=\begin{pmatrix}0.08&-0.19&0.20\\ -0.23&0.31&-0.23\\ 0.18&-0.17&-0.16\end{pmatrix},\qquad P_{2}^{\prime}=\begin{pmatrix}-0.31&0.03&0.11\\ 0.06&-0.06&0.13\\ 0.14&0.11&0.10\end{pmatrix}$.
The matrices $D_{1}(t)$ and $D_{2}(t)$ were given by:
$D_{1}(t)$ $=2\operatorname{diag}\big(\cos(10\pi t),~\sin(10\pi t),~\cos(6\pi t)\big),$
$D_{2}(t)$ $=2\operatorname{diag}\big(\cos(6\pi t),~\sin(6\pi t),~\cos(4\pi t)\big),$
where $\operatorname{diag}:\mathbb{R}^{3}\to\mathcal{M}_{3\times 3}(\mathbb{R})$ denotes the function that maps a vector to the diagonal matrix with its components on the diagonal.
Figure 1. Frobenius norm of the matrices $P_{1}(t)$ and $P_{2}(t)$.
@@
Figure 2. Convergence to a consensus equilibrium on the sphere $\mathbb{S}^{2}$. All the tokens start and remain in an hemisphere.
To better appreciate the time-varying nature of the matrices $P_{1}$ and $P_{2}$, in Fig.˜1 we shown their Frobenius norm.
In Fig.˜2 we show the motion of the tokens in blue with their initial position represented by a white circle and final position by a gray circle. We can appreciate that all the tokens start and remain in an hemisphere and that they converge to a consensus equilibrium.
@@
The proof of Theorem˜4.1 is based on the Lyapunov function Eq.˜10, whose time-evolution is displayed in Fig.˜3 for the case where $v=(1,0,0)$.
Figure 3. Evolution of the Lyapunov function Eq.˜10 used in the proof of Theorem˜4.1.
6.1.2. Illustration of Theorem˜5.1
@@
We now consider the auto-regressive model with $50$ tokens on $\mathbb{S}^{499}\subset\mathbb{R}^{500}$. The number and dimension of the tokens were chosen to make them comparable to the GPT-2 model. We use two heads ($h=2$) with the matrices $P_{1}=D_{1}(t)P_{1}^{\prime}$ and $P_{2}=D_{2}(t)P_{2}^{\prime}$ obtained by randomly generating $P_{1}^{\prime}$ and $P_{2}^{\prime}$, and taking $D_{1}(t)$ and $D_{2}(t)$ to be diagonal with entries $(D_{\eta})_{jj}=|2\sin(wt+\phi)|$ for $\eta=1,2$, $j=1,\ldots,500$, $w$ drawn from the uniform distribution on ]0,1[]0,1[ and $\phi$ drawn from the uniform distribution on $0,2\pi[$. To measure the error between tokens we use the cosine similarity, $E:(\mathcal{E}_{W}^{n})^{\ell}\to\mathbb{R}^{+}$, defined as:
(25) $E=1-\frac{1}{\ell}\sum_{i=1}^{\ell}\frac{{\rm y}_{1}^{\top}{\rm y}_{i}}{|{\rm y}_{1}|\,|{\rm y}_{i}|},$
which becomes zero when all the tokens belong to the consensus set.
@@
In Fig.˜4 we display the evolution of the function $E$ along $100$ trajectories of Eq.˜6 for random initial conditions drawn from an element-wise uniform distribution on ]0.5,0.5[]-0.5,0.5[, and then projected to the sphere. We can appreciate in Fig.˜4 how the function $E$ converges to zero along all the trajectories.
@@
Figure 4. Illustration of Theorem˜5.1; evolution of the function $E$ defined in Eq.˜25 along $100$ solutions of Eq.˜6 with random initial conditions drawn from an element-wise uniform distribution on ]0.5,0.5[]-0.5,0.5[ and then projected to the sphere.
6.1.3. Illustration of Theorem˜5.2
In the final case we use the causal model with $10$ tokens on $\mathbb{S}^{2}$ with randomly assigned initial positions. As for the previous cases, we choose $P(t)=D(t)P^{\prime}$, with randomly generated $P^{\prime}$ and $U$ given by:
$P^{\prime}=\begin{pmatrix}0.36&0.42&0.13\\ 0.10&-0.07&-0.20\\ 0.15&-0.21&0.12\end{pmatrix},\qquad U=\begin{pmatrix}-0.26&0.50&0.56\\ 0.50&-0.72&-0.50\\ 0.56&-0.50&-0.02\end{pmatrix}$,
and $D(t)=2\operatorname{diag}\big(\cos(10\pi t),~\sin(10\pi t),~\cos(6\pi t)\big)$.
In Fig.˜5 we can observe convergence of the tokens to a consensus equilibrium point whereas in Fig.˜6 we have the time evolution of $V_{1}=1-{\rm y}_{1}^{\top}v$ and $V_{2}=1-{\rm y}_{2}^{\top}v$ where $v\in\mathbb{R}^{3}$ is the eigenvector of $U$ corresponding to its largest eigenvalue. Note that $V_{2}$ is not a Lyapunov function, and therefore it may increase, although the proof of Theorem˜5.2, establishes that it will eventually converge to zero.
Figure 5. Convergence to a consensus equilibrium on the sphere $\mathbb{S}^{2}$. All the tokens start and remain in the hemisphere defined by $v$.
Figure 6. Evolution of the functions $V_{1}$ and $V_{2}$ used in the proof of Theorem˜5.2.
6.2. GPT-2 and GPT-Neo Experiments